Helium gas can be used to displace other gases from a solvent by bubbling helium through the solvent, a process called sparging, and leaving an atmosphere of helium above the solvent. At 760.0 torr of helium, the concentration of helium in water is 0.00148 g L^-1 at 298 K. What is Henry’s law constant for helium at 298 K in g L^-1 torr ^-1?
Answer #1

Given that, Pressure = 760 torr

Concentration = 0.00148 g/L

From Henry's Law which states that, the amount of gas dissolved in a given volume of liquid is directly proportional to pressure applied

Mathematically,

Concentration dissolved Pressure

Concentration = KH * Pressure

Where KH is the henry's constant

So, 0.00148 g/L = KH * 760 torr

KH = 1.95 *10-6 g/L.torr